# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

### v2

@@ 6 Discussion
We find that despite the many experiments we ran, including ones with different configurations, failure modes were extremely rare. In fact, we didn’t experience any failure modes except with the gte model. Moreover, all failure modes were around the gte-gtr pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in gtr $\rightarrow$ gte – they are just rarer. Crucially, failure modes are not catastrophic. The failure modes are characterized by worse top-1 accuracy and average rank (and slightly lower cosine similarity throughout the refinement steps. Often also noticeable in the initial step, but not always). But the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
@@ 7 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. Lenc and Vedaldi (2015) and Bansal et al. (2021) demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. Olah (2015) studied the equivalence of neural networks in relative terms. Moschella et al. (2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Similarly, Dar et al. (2023) have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors. Huh et al. (2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of unsupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018). Most of the works rely on training a GAN. They often require a small seed of matched pairs in order to work. Hoshen and Wolf (2018) avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, specifically their approximate matching process, which relied on PCA, was insufficient for our purposes.
@@ 8 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, provided that structural correspondences can be estimated accurately.
The success of our approach provides strong empirical evidence for the universal geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
The efficiency and accessibility of our method could democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al. (2025).

### v3

@@ 6.2 Failure Modes
We find that despite the many experiments we ran, ranging diverse hyperparameter configurations, failure modes were extremely rare. In fact, we didn’t experience any except on the $gte\leftrightarrow gtr$ pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in the gtr $\rightarrow$ gte direction, they are just rarer. Crucially, failure modes are not catastrophic. Failure modes are characterized by worse top-1 accuracy and average rank – but the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
6.3 Alternative Approaches
Through anecdotal experiments, we have found that it might be possible to sidestep the use of relative representations in the approximate matching step. Instead, we can perform Procrustes Analysis directly on the concatenated aligned centroids. It seems to do slightly worse in the approximate matching step, but it converges normally after. We emphasize, as noted above, that we haven’t tested this thoroughly to be certain it generalizes. For future reference, we call this variant mini-vec2vec without relative representations”, but catchier names can be used if meaning is clear.
@@ 7 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. lenc2014stitching and bansal2021revisiting have demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. colah2015visualizing studied the equivalence of neural networks in relative terms. moschella2022relative have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Similarly, dar2023analyzing have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors. huh2024platonic have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of unsupervised alignment of embeddings also has a long history in word embeddings (conneau2017word; chen-cardie-2018-unsupervised; grave2018unsupervisedalignmentembeddingswasserstein) and machine translation (lample2018unsupervisedmachinetranslationusing; artetxe2018unsupervisedneuralmachinetranslation). Most of the methods rely on training a GAN, and they often require a small seed of matched pairs in order to work. hoshen-wolf-2018-non avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, especially the approximate matching step, which relied on PCA, was insufficient for our purposes.
@@ 8 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways. The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in jha2025vec2vec.